C=1500+.6 (YD)

I=2400-10r

T=1800+.2t

G=2000

NX=-200

L=.2y -50r

M/P=1000

A. Compute the value of the multiplier. B. Compute the value of the autonomous planned spending. C. Write an equation representing the "IS" curve. D. Write an equation representing the "LM" curve. E. Find equilibrium interest rate and real GDP. F. Calculate the crowding effect of additional $500 government expenditures.



1
Expert's answer
2021-07-26T17:01:04-0400

A. The value of the multiplier is:

"m = 1\/(1 - c) = 1\/(1 - 0.6) = 2.5."

B. The value of the autonomous planned spending is:

Sa = Ca + Ia + Ga + NXa = 1,500 + 2,400 + 2,000 - 200 = 5,700.

C. An equation representing the "IS" curve is:

Y = C + I + G + NX = 1,500 + 0.6(Y - 1,800 - 0.2r) + 2,400 - 10r + 2,000 - 200 = 0.6Y + 4,620 - 10.12r,

0.4Y = 4,620 - 10.12r,

Y = 11,550 - 25.3r

D. An equation representing the "LM" curve is:

L = 0.2y - 50r = M/P = 1,000,

Y = 5,000 + 250r,

E. The equilibrium interest rate and real GDP are:

11,550 - 25.3r = 5,000 + 250r,

275.3r = 6,550,

r = 23.79%,

Y = 5,000 + 250×23.79 = 10,947.5.

F. The crowding effect of additional $500 government expenditures means, that the real GDP may increase less than is planned according to the multiplier effect.